OpenStudy (anonymous):

A cyclist rode 3 hours with a tail of 6 km/h. The return trip against the head wind took 5 hours. Find the cyclist's speed in still air

OpenStudy (nurali):

3(x+6)=5(x-6) 3x+18=5x-30 -2x=-48 Dividing by 2 x=?